# Register Any Point: Scaling 3D Point Cloud Registration by Flow Matching

### v1

@@ 30       SaveTrainingSample($\mathcal{M}$);
where TE is measured in meters and RE is measured in degrees. For multi-view registration, we report the mean translation error and mean rotation error across all point clouds.
For multi-view registration, we also evaluate the quality of the registered point cloud using the Chamfer Distance (CD), which measures the bi-directional root mean squared distance between the registered point cloud and the ground truth aggregated point cloud. Given the registered point cloud $\mathbf{P}_{\text{reg}}$ and the ground truth point cloud $\mathbf{P}_{\text{gt}}$, the Chamfer Distance is computed as:
$\text{CD}=\sqrt{\frac{1}{2}\left(\frac{1}{|\mathbf{P}_{\text{reg}}|}\sum_{\mathbf{p}\in\mathbf{P}_{\text{reg}}}\min_{\mathbf{q}\in\mathbf{P}_{\text{gt}}}\|\mathbf{p}-\mathbf{q}\|_{2}^{2}+\frac{1}{|\mathbf{P}_{\text{gt}}|}\sum_{\mathbf{q}\in\mathbf{P}_{\text{gt}}}\min_{\mathbf{p}\in\mathbf{P}_{\text{reg}}}\|\mathbf{q}-\mathbf{p}\|_{2}^{2}\right)},$
where CD is measured in meters.
Appendix B Details on the Selected Baselines
In the main paper, we did not include detailed descriptions of the baseline methods due to space limitations. Here we provide comprehensive descriptions of all baselines used in our experiments in Tab. 7.
All pairwise registration baselines follow the standard correspondence matching and transformation estimation pipeline. Similarly, all multi-view registration baselines follow a two-stage pipeline: first performing pairwise registration, then applying pose graph optimization to enforce global consistency. In contrast, our method is a single-stage multi-view registration approach that directly generates the registered point cloud using flow matching, eliminating the need for exhaustive pairwise correspondence matching, pairwise transformation estimation, and pose graph optimization.
@@ Appendix B Details on the Selected Baselines
Pairwise Registration Baselines (all follow the correspondence matching and transformation estimation pipeline)
@@
FPFH [rusu2009icra] + FGR [zhou2016eccv] Conventional ECCV’16 Handcrafted Fast Point Feature Histogram (FPFH) descriptor extracts local geometric features for correspondence matching, combined with Fast Global Registration (FGR) that formulates registration as a global optimization problem using a robust cost function to handle outliers effectively
@@
FPFH [rusu2009icra] + TEASER [yang2020tro] Conventional TRO’20 Handcrafted FPFH descriptor for correspondence matching combined with TEASER (Truncated least squares Estimation And SEmidefinite Relaxation), which uses truncated least squares and semidefinite relaxation to provide certifiably optimal solutions robust to extreme outlier rates
FPFH [rusu2009icra] + Quatro [lim2022icra] Conventional ICRA’22 Handcrafted FPFH descriptor for correspondence matching combined with Quatro, a Quasi-SE(3) registration method that extends TEASER to handle degenerate cases in urban environments by requiring only a single correspondence
@@
FCGF [choy2019iccv] Deep learning ICCV’19 Fully Convolutional Geometric Features uses a sparse 3D convolutional network to extract dense geometric features from point clouds, enabling dense correspondence matching for robust registration
@@
Predator [huang2021cvpr-predator] Deep learning CVPR’21 Overlap-aware registration network with attention mechanism that learns to focus on overlapping regions between point clouds, particularly effective for low-overlap scenarios by predicting overlap scores and using attention-weighted features for correspondence matching and transformation estimation
GeoTransformer [qin2022cvpr-geotransformer] Deep learning CVPR’22 Keypoint-free registration method that matches superpoints based on patch overlap. It encodes pair-wise distances and triplet-wise angles to learn transformation-invariant geometric features through self-attention, enabling robust superpoint matching in low-overlap scenarios. Dense correspondences are obtained via optimal transport, achieving high inlier ratios without requiring RANSAC
BUFFER [ao2023cvpr-buffer] Deep learning CVPR’23 Registration method that balances accuracy, efficiency, and generalizability through three key components: a Point-wise Learner that predicts keypoints and estimates point orientations, a Patch-wise Embedder that extracts efficient and general patch features, and an Inliers Generator that identifies inlier correspondences for robust transformation estimation
EYOC [liu2024cvpr-eyoc] Deep learning CVPR’24 Unsupervised method for distant point cloud registration that progressively trains a feature extractor using near point cloud pairs to self-supervise on farther pairs, eliminating the need for global pose labels through progressive distance extension, works well for outdoor low-overlap scenarios

### v2

@@ B.6 Evaluation metrics
where TE is in meters and RE is in degrees. For multi-view registration, we report the mean TE and RE across all scans in a sample.
For multi-view registration on the cross-domain benchmark, we adopt the edge success rate and graph success rate as the primary evaluation metrics. The edge success rate is the mean registration success rate over all valid edges (i.e., overlapping pairs) in a sample. The graph success rate is stricter: it requires every valid edge in the multi-view registration graph to be successful, thus measuring the overall consistency of the registration. To handle diverse scene scales, we normalize the translation error by dividing by the longest axis of the bounding box of the ground-truth registered point cloud; a registration is considered successful if the normalized translation error is below 2.5% and the rotation error is below 15°. We also report results under stricter thresholds (0.5%, 3°) in this supplementary material.
For the additional multi-view registration experiments in this supplementary material, we adopt the following metrics: Chamfer distance (CD), normalized global RMSE, and the empirical cumulative distribution function (ECDF) of error. The CD measures the bi-directional root-mean-squared distance between the registered point cloud and the ground-truth aggregated point cloud. Given the registered point cloud $\mathbf{P}_{\text{reg}}$ and the ground-truth point cloud $\mathbf{P}_{\text{gt}}$, the CD is computed as:
$\small\text{CD}=\sqrt{\frac{1}{2}\left(\frac{1}{|\mathbf{P}_{\text{reg}}|}\sum_{\mathbf{p}\in\mathbf{P}_{\text{reg}}}\min_{\mathbf{q}\in\mathbf{P}_{\text{gt}}}\|\mathbf{p}-\mathbf{q}\|_{2}^{2}+\frac{1}{|\mathbf{P}_{\text{gt}}|}\sum_{\mathbf{q}\in\mathbf{P}_{\text{gt}}}\min_{\mathbf{p}\in\mathbf{P}_{\text{reg}}}\|\mathbf{q}-\mathbf{p}\|_{2}^{2}\right)}.$ (9)
The global RMSE is computed as the root-mean-squared error between the generated point cloud and the ground-truth registered point cloud using the known point-to-point correspondences, and is normalized by the longest axis of the ground-truth registered point cloud’s bounding box to account for diverse scene scales.
The ECDF of error reports the fraction of samples whose error falls below a given threshold. Plotting the ECDF across thresholds summarizes the full error distribution and allows comparison of methods at different accuracy levels.
Appendix C Details on the Selected Baselines
In the main paper, we did not include detailed descriptions of the baseline methods due to space limitations. Here we provide comprehensive descriptions and used training datasets of all baselines used in our experiments in Tab.˜6.
All pairwise registration baselines follow the standard correspondence matching and transformation estimation pipeline. All multi-view registration baselines except RPF [sun2025neurips] follow a two-stage pipeline: first performing pairwise registration along dense or sparse graph edges, then applying pose graph optimization to enforce global consistency. In contrast, our method is single-stage and directly generates the registered point cloud via flow matching, eliminating the need for exhaustive pairwise correspondence matching, transformation estimation, and pose graph optimization.
When evaluating on the cross-domain benchmark, we select models for learning-based baselines as follows. For methods with multiple models trained on different datasets, we choose based on scene scale: ModelNet for object-centric, 3DMatch for indoor small-scale, and KITTI for outdoor large-scale scenes. For methods with only a single model, we rescale the input to match the training scale at test time and then scale back for evaluation.
@@ Appendix C Details on the Selected Baselines
Pairwise Registration Baselines
@@
FPFH [rusu2009icra] + FGR [zhou2016eccv] Conventional ECCV’16 Handcrafted FPFH features combined with fast global registration via a robust cost function.
@@
FPFH [rusu2009icra] + TEASER [yang2020tro] Conventional TRO’20 Handcrafted FPFH features with certifiably optimal pose estimation via truncated least squares and semidefinite relaxation, robust to high outlier rates.
KISS-Matcher [lim2025icra-kissmatcher] Conventional ICRA’25 Fast and robust registration combining efficient hand-crafted feature extraction with robust correspondence estimation, designed for scalability and generalization across diverse scenes.
@@
FCGF [choy2019iccv] Deep learning ICCV’19 Sparse 3D convolutional network for dense geometric feature extraction, enabling dense correspondence matching for robust registration. Trained on 3DMatch or KITTI.
@@
Predator [huang2021cvpr-predator] Deep learning CVPR’21 Overlap-aware network that predicts overlap scores and uses attention-weighted features to focus on overlapping regions, particularly effective in low-overlap scenarios. Trained on 3DMatch, KITTI or ModelNet.
GeoTransformer [qin2022cvpr-geotransformer] Deep learning CVPR’22 Keypoint-free method matching superpoints via transformation-invariant geometric features; uses optimal transport for dense correspondences without RANSAC. Trained on 3DMatch, KITTI or ModelNet.
BUFFER [ao2023cvpr-buffer] Deep learning CVPR’23 Balances accuracy and efficiency via learned keypoint detection, patch feature embedding, and inlier correspondence generation. Trained on 3DMatch or KITTI.